Complete {(2,0,3)} to form an orthogonal basis of R^3


1
Expert's answer
2021-02-24T12:24:08-0500

Solution:

According to the Gram-Schmidt process,

"\mathbf{u}_{\mathbf{k}}=\mathbf{v}_{\mathbf{k}}-\sum_{j=1}^{k-1} \operatorname{proj}_{\mathbf{u}_{\mathbf{j}}}\left(\mathbf{v}_{\mathbf{k}}\right)" , where

"\operatorname{proj}_{\mathbf{u}}(\mathbf{v})=\dfrac{\mathbf{u} \cdot \mathbf{v}}{\mathbf{u} \cdot \mathbf{u}} \mathbf{u}"

The normalized vector is 

"\mathbf{e}_{\mathbf{k}}=\dfrac{\mathbf{u}_{\mathbf{k}}}{\sqrt{\mathbf{u}_{\mathbf{k}} \cdot \mathbf{u}_{\mathbf{k}}}}"

Here "\cdot" is the dot product operator

"\mathbf{u}_{\mathbf{1}}=\mathbf{v}_{\mathbf{1}}=\left[\begin{array}{l}2 \\ 0 \\ 3\end{array}\right]"

Now, "\mathbf{e}_{1}=\dfrac{\mathbf{u}_{1}}{\sqrt{\mathbf{u}_{1} \cdot \mathbf{u}_{1}}}"

"\sqrt{\mathbf{u}_{1} \cdot \mathbf{u}_{1}}=\sqrt{2^2+0^2+3^2}=\sqrt{4+9}=\sqrt{13}"

So, "\mathbf{e}_{1}=\dfrac{\mathbf{u}_{1}}{\sqrt{\mathbf{u}_{1} \cdot \mathbf{u}_{1}}}=\left[\begin{array}{c}\dfrac{2 \sqrt{13}}{13} \\ 0 \\ \dfrac{3 \sqrt{13}}{13}\end{array}\right]"

So, the required set of orthonormal vectors is "\mathbf{e}_{1}=\left[\begin{array}{c}\dfrac{2 \sqrt{13}}{13} \\ 0 \\ \dfrac{3 \sqrt{13}}{13}\end{array}\right]"